OpenStudy (anonymous):

Find the length of the hypotenuse of a right triangle whose legs are 5 and sqrt 2

OpenStudy (anonymous):

(5)^2+(sqrt2)^2=c^2 25+2=c^2 27=c^2 sqrt27=c 3sqrt3=c